# Generative Logic: A New Computer Architecture for Deterministic Reasoning and Knowledge Generation

### v2

@@ 6.4 LLM Deployment
LLM Interface & Feedback Loop (post‑ENT). After the ENT milestone, we will expose tool calls so an LLM can (i) submit MPL definitions/conjectures, (ii) invoke GL runs, and (iii) inspect derived theorems and proof graphs. In the reverse direction, GL becomes an LLM reasoning booster: when the model confronts a mathematical (and eventually any MPL‑encodable) reasoning task, it can translate that task to MPL and ask GL to deterministically study its deductive consequences, feeding validated results back into the dialogue. GL remains an external proof engine—not a Python library import—so the LLM/GL handshake mirrors, but does not collapse into, standard Python execution flows.
@@ 7 Potential Impact (Speculative)
7.1 LLMs Integration
Bidirectional LLM–GL Synergy. Today’s LLM coding assistants improve developer productivity by writing Python that a trusted interpreter executes. We aim for an analogous—but higher‑stakes—loop in formal reasoning: LLMs learn the MPL used by GL, submit definitional fragments or conjectures, and invoke GL’s deterministic, definition‑first exploration engine to mine consequences and produce fully traceable proof graphs. In return, GL functions as an LLM enhancer: whenever the model faces a mathematical (and potentially broader structured) reasoning task, it can externalize that task into MPL, let GL study it rigorously, and incorporate the resulting validated theorems—or contradictions—back into its natural‑language deliberation. Because GL records provenance for every derived step and was designed for scalable, auditable theorem generation from definitions, this closed loop could add a layer of trust and rigor to LLM‑mediated mathematical workflows well beyond what probabilistic text generation alone can provide. (Exploratory; not yet implemented.)
@@ 8 Conclusion
Generative Logic (GL) set out to test a simple question: Can a deterministic, definition‑centric architecture automatically discover and verify non‑trivial theorems from first principles? In this paper we demonstrated an end‑to‑end baseline: starting from a formalised Peano axiom set in MPL, GL generated conjectures and constructed verifiable proofs for the six foundational arithmetic laws, emitting a fully navigable HTML proof graph for audit.
Three architectural properties were validated empirically. First, the hash‑table inference model enables transparent, stepwise replay of every derivation. Second, compilation of definitions into distributed logic blocks cleanly separates symbolic content from execution, supporting parallel scaling. Third, the definition‑first workflow shifts human effort to axiom design; once definitions compile, proofs follow automatically (Section 2).
The current implementation remains a point‑zero prototype: single‑core Python, string‑based data structures, and modest performance headroom. Section 6 outlines the path to scale—C++ realisation, richer theorem weaving (negation, trichotomy)—on the way to ENT and beyond.
We view GL not as a replacement for human mathematicians but as a new instrument: a deterministic engine that can explore the deductive consequences of axioms at machine speed, extending the reach of formal reasoning.
These directions are intentionally forward‑looking; we present the Peano baseline in this paper to enable the community to evaluate, replicate, and extend the approach.

### v3

@@ 6.7 LLM Deployment
LLM Interface & Feedback Loop (post-ENT). After the ENT milestone, we will expose tool calls so an LLM can (i) submit MPL definitions and conjectures, (ii) invoke GL runs, and (iii) inspect derived theorems and proof graphs. In the reverse direction, GL becomes an LLM reasoning booster: when the model confronts a mathematical (and eventually any MPL-encodable) reasoning task, it can translate that task to MPL and ask GL to deterministically study its deductive consequences, feeding validated results back into the dialogue. GL remains an external proof engine — not a Python library import — so the LLM/GL handshake mirrors, but does not collapse into, standard Python execution flows.
@@ 8 Conclusion
GL set out to answer a narrow question: whether a deterministic, definition-centric architecture can autonomously generate and verify non-trivial theorems from first principles, without human guidance at the level of individual proof steps.
The question has been answered in the affirmative, at least for the domain examined. Starting from a formalized set of Peano axioms, the system derived the foundational laws of arithmetic. Starting from those results, it derived Gauss’s summation formula. Both derivations are recorded as complete proof graphs in which every inference step is explicitly linked to its antecedent, enabling independent verification by navigating or scripting over the generated output.
Three properties of the architecture were confirmed in practice. First, every inference step is a hash-table lookup, which means every derivation is replayable and independently checkable — the proof graph is not a summary but a complete record. Second, each Logic Block holds only its local fragment of the reasoning task; blocks communicate between cycles, not during them. This clean separation is what makes the architecture parallelizable in principle and, eventually, on hardware. Third, the system was given definitions. It was not given theorems. The theorems it produced were not requested. This is the property the architecture was designed around, and it held.
The implementation is a prototype. Its performance figures are those of a prototype. The path to larger theories is identified and involves no steps that appear, at present, to be impossible.
Whether the architecture scales to the FTA and beyond remains to be established. We expect that it does.